OpenStudy (anonymous):

Determine the slope of the line containing these two points. Please show all of your work. Then describe what a line with this slope would look like. (-1, -5) and (-3, 1)

OpenStudy (anonymous):

the line has slope of -3, and so you can work out the equation y + 5 = -3x + 3 y = -3x -2

OpenStudy (hoblos):

slope = (y2-y1)/(x2-x1) = (1-(-5))/(-3-(-1)) = 6/-2 = -3

OpenStudy (anonymous):

so it is quite steep, and looks like below

OpenStudy (anonymous):

sorry, the line should be lower, can't draw very well